If 376 mL of 5.52 magnesium sulfide is reacted with excess aluminum chloride according to the following balance equation
3MgS + 2AlCL3— 3MgCl2 + Al2S3
1
Expert's answer
2019-02-20T04:39:19-0500

3MgS + 2AlCl3 -> 3MgCl2 + Al2S3

n=m/M

M(MgS) = 24.31+32.07 = 56.38 g/mol

n(MgS) = 5.52 g/ 56.38 g/mol = 0.0979 mol

According to equation 3 mol of MgS give 1 mole of Al2S3

We have 0.0979 mol of MgS which give x mol of Al2S3

Solve the proportion:

3/0.0979 = 1/x

x = 0.0326

n(Al2S3)= 0.0326 mol

m = M*n

M(Al2S3) = 26.98*2+32.07*3 = 150.17 g/mol

m(Al2S3) = 0.0326 mol *150.17 g/mol = 4.90 g

Answer: 4.90 g